Consider the following simplified financial statements for the Steveston Corporation (assuming no income taxes):


Statement of Comprehensive Income Statement of Financial Position
  Sales $ 32,000   Assets $ 25,300   Debt $ 5,800
  Costs 24,400   Equity 19,500
    Net income $ 7,600     Total $ 25,300     Total $ 25,300


Steveston has predicted a sales increase of 15 percent. Assume Steveston pays out half of net income in the form of a cash dividend. Costs and assets vary with sales, but debt and equity do not.

  

Prepare the pro forma statements. (Input all amounts as positive values.)


Pro forma Statement of Comprehensive Income Pro forma Statement of Financial Position
  Sales $ Assets $ Debt $
  Costs Equity
  Net income $ Total $ Total $


Determine the external financing needed. (Negative amount should be indicated by a minus sign.)


  External financing needed $   
Answer #1

Answer

Pro-forma Income Statement:

Sales = 32000 x (1+15%) = 36,800

Cost = 24000 x (1+15%) = 27,000

Pro-forma Balance Sheet without Considering Dividend:

Assets = 25,300 x (1+15%) = 29095

Cash is the Plug figure = 34040 - 29095 = 4945

Half of the Net Income is paid out in Cash so

Retained Earning = 8740/2 =4370

Proforma Balance Sheet:

Here too no EFN is needed. We need a plug figure on asset side which will again be cash

Cash Plug = 29670 - 29095 = 575

Please private rating